If the caffeine concentration in a particular brand of soda is
1.71
mg/oz,
drinking how many cans of soda would be lethal? Assume that 10.0 g of caffeine is a lethal dose, and there are 12 oz in a can.
1
Expert's answer
2020-02-18T06:37:30-0500

Amount of caffeine present in one can = 1.71*12 = 20.52 mg

10.0 g of caffeine is a lethal dose

So number of cans that will constitute to 10 grams of caffeine = "\dfrac{10}{20.52\times10^{-3}}"= "487" cans